OpenStudy (anonymous):

What is the volume of a cylinder with a height of 7 cm and a radius of 9 cm? plz help

OpenStudy (kamibug):

Here is the formula for the volume of a cylinder. :)

OpenStudy (kamibug):

Plug in the height and radius and solve... V = pi*9^2*7 Can you solve that? :D

OpenStudy (anonymous):

no i suck at math can you explain how i would solve that plz

OpenStudy (kamibug):

No problem. ^.^ First find what 9 squared is. What's 9*9?

OpenStudy (anonymous):

81

OpenStudy (kamibug):

Right! So now we have: V = pi*81*7 Let's multiply 81 by pi (3.14) now. What's 81* 3.14? :)

OpenStudy (anonymous):

how did you get 3.14 ??

OpenStudy (kamibug):

3.14 is pi. That's what the sign \[\pi \] in the formula means. :)

OpenStudy (anonymous):

oh that makes sense so now i need to multiply 81*3.14 ??

OpenStudy (kamibug):

Yup! :3

OpenStudy (anonymous):

ok i get 254.34 right or wrong??

OpenStudy (kamibug):

That's right. :D Now the last step is to multiply that by the height, 7. 254.34 * 7 = ?

OpenStudy (anonymous):

1780.38 ??

OpenStudy (kamibug):

That's correct. :) Now you have the volume. :D Great job! ^.^